### 10 Mar 2026 v1

# The Epistemic Support-Point Filter: Jaynesian Maximum Entropy Meets Popperian Falsification A Possibilistic Minimax-Entropy Optimality Proof

Abstract
The Epistemic Support-Point Filter (ESPF) was designed around a single epistemological commitment: be quick to embrace ignorance and slow to assert certainty. This paper proves that this commitment has a precise mathematical form and that the ESPF is the unique optimal filter implementing it within the class of epistemically admissible evidence-only filters.
The ESPF operates as the synthesis of two complementary epistemological principles acting at different phases of the recursion. In its propagation phase it enacts Jaynesian maximum entropy [15]: the support spreads as widely as the dynamics and process noise allow, assuming maximal ignorance consistent with the known dynamical constraints. In its measurement-update phase it enacts Popperian falsification [16]: hypotheses are eliminated if and only if the evidence rules them out, using evidence compatibility alone. Any selection rule that incorporates prior possibility is strictly suboptimal and risks systematic race-to-bottom bias over multiple steps. The filter alternates between these two operations at every step: expand maximally, then contract minimally.
The optimality criterion is possibilistic minimax entropy: among all evidence-only survivor selection rules, the ESPF’s minimum-$q$ selection minimizes the maximum possibilistic entropy consistent with the surviving hypotheses—equivalently, it minimizes $\log\det(\mathrm{MVEE})$, the log-volume of the outermost $\alpha$-cut.
Three lemmas establish the result. The Possibilistic Entropy Lemma identifies the possibilistic entropy $H_{\pi}$—the integral of $\log V_{\alpha}$ over $\alpha\in(0,1]$—as the correct ignorance functional and decomposes it into a minimax term and a gradient term. The Possibilistic Cramér–Rao Lemma bounds how rapidly any admissible filter may reduce $H_{\pi}$ per measurement. The Evidence-Optimality Lemma proves that evidence-only selection minimizes $\log\det(\mathrm{MVEE})$ at every $\alpha$-level simultaneously.
Numerical validation over a 2-day, 877-step Level-3 Smolyak run ($M=106$ support points, $n=7$) under both nominal LEO tracking and a stress case (10 m/s crosstrack maneuver ++ 20 m Arecibo range bias) confirms the theorem’s regime structure. The Epistemic Width Monitor (EWM) diagnostic suite reveals that possibilistic stress under sustained model mismatch manifests primarily through necessity saturation and surprisal escalation rather than MVEE sign change, establishing these as the more sensitive early-warning indicators of model–reality divergence. The Kalman filter is recovered in the Gaussian limit.
Keywords: Possibility theory; Possibilistic entropy; Minimax entropy; $\alpha$-cut; Popperian falsification; Jaynesian maximum entropy; Possibilistic Cramér–Rao bound; Non-Bayesian estimation; Kalman filter generalization; Epistemic Width Monitor.
1 Introduction
1.1 Two Principles, One Filter
There is a tension at the heart of recursive state estimation. On one side stands E. T. Jaynes [15]: when you do not know something, assume the maximum entropy consistent with what you do know. Assert nothing you have not earned. On the other side stands Karl Popper [16]: when evidence arrives, eliminate every hypothesis the evidence rules out. Keep only what the data has not falsified.
These two principles appear to pull in opposite directions. Jaynes says spread; Popper says cut. Yet they are not in conflict. They operate at different temporal phases of a single recursive epistemology. Before a measurement, Jaynes governs: propagate the possibility distribution forward under maximal ignorance. After a measurement, Popper governs: eliminate hypotheses incompatible with the evidence, using the evidence alone. The filter that implements this two-phase epistemology is the ESPF.
The design intent can be stated simply: be quick to embrace ignorance and slow to assert certainty. This paper proves that this intent is a theorem.
The ESPF differs from Bayesian filters in a foundational way: it minimizes worst-case epistemic ignorance rather than expected uncertainty. A Bayesian filter integrates over a prior to produce a posterior; the ESPF selects survivors by evidence alone and minimizes the maximum residual ignorance consistent with those survivors. Where prior information is unreliable or unavailable, this distinction is not stylistic—it is the difference between an epistemically defensible recursion and one that risks compounding prior errors.
1.2 The Minimax Entropy Principle
The key to making the Popperian update optimal is identifying the right criterion. The right criterion is not minimizing expected entropy, or minimizing posterior-weighted entropy, or any criterion that mixes evidence with prior belief. The right criterion is minimax entropy: among all evidence-only selection rules, choose the one that minimizes the worst-case residual ignorance.
The worst-case ignorance after retaining a survivor set $J$ is $\log\det(\mathrm{MVEE}(\{\chi^{(i)}:i\in J\}))$—the log-volume of the outermost $\alpha$-cut of the surviving possibility distribution. This is the entropy of the most pessimistic agent who considers all survivors equally plausible. Minimizing it yields the set of $N_{\mathrm{target}}$ survivors whose convex hull is as small as possible: the hypotheses with smallest whitened squared innovations $q^{(i)}_{k}$, geometrically closest to the evidence-consistent locus. This is exactly the ESPF’s selection rule.
1.3 The Two Regimes
The Jaynesian-Popperian synthesis has a natural regime structure.
Falsification regime ($\log\det(\mathrm{MVEE})<0$): the possibility support occupies a volume smaller than the unit sphere in innovation-normalized coordinates. The Popperian principle is operative and the minimax entropy theorem applies.
Diffusion regime ($\log\det(\mathrm{MVEE})\geq 0$): the support has spread beyond the unit sphere due to process noise accumulation. The filter maintains coverage of the reachable state space; the minimax entropy criterion does not apply. This is the Jaynesian phase, directly analogous to the Kalman filter’s process-noise prediction step during which the measurement-optimality criterion is inoperative.
The regime boundary at $\log\det(\mathrm{MVEE})=0$ is a real-time diagnostic indicator of which epistemological principle governs the current step.
1.4 The Correct Entropy Functional
For a uniform possibility distribution, log-volume suffices. But the ESPF’s distribution is generically non-uniform after the first update: the conjunctive update assigns each survivor $\tilde{\pi}^{(i)}_{k}=\min(\pi^{(i)}_{k|k-1},\mathrm{Comp}^{(i)}_{k})$, varying across survivors according to their evidence compatibility. The correct functional is the possibilistic entropy $H_{\pi}$, defined as the integral of $\log V_{\alpha}$ over all $\alpha$-levels. It decomposes into a minimax term (the $\alpha\to 0$ limit, recovering MVEE log-det) and a gradient term (negative, measuring how sharply possibility is concentrated within the admissible set). The ESPF minimizes both simultaneously.
1.5 Structure of the Paper
Section 2 establishes the framework. Section 3 proves the three core lemmas. Section 4 states and proves the main optimality theorem. Section 5 establishes the Kalman filter as the Gaussian limit. Section 6 presents numerical validation at Smolyak Level 3 over a 2-day orbital tracking run including the full EWM diagnostic suite. Section 7 discusses implications.
2 Setup
2.1 Possibility Distributions, $\alpha$-Cuts, and Possibilistic Entropy
Definition 2.1 (Possibility distribution and $\alpha$-cuts).
A possibility distribution over a finite support $\{\chi^{(i)}\}_{i=1}^{M}\subset\mathbb{R}^{n}$ is a function $\pi:\{1,\ldots,M\}\to(0,1]$ with $\max_{i}\pi^{(i)}=1$. The $\alpha$-cut at level $\alpha\in(0,1]$ is
$C_{\alpha}=\{i:\pi^{(i)}\geq\alpha\}.$ (1)
The $\alpha$-cuts are nested: $C_{\alpha^{\prime}}\subseteq C_{\alpha}$ for $\alpha^{\prime}>\alpha$. The volume of the $\alpha$-cut is $V_{\alpha}=c_{n}(\det\Pi_{\alpha})^{1/2}$, where $\Pi_{\alpha}\succ 0$ is the MVEE shape matrix of $\{\chi^{(i)}:i\in C_{\alpha}\}$ and $c_{n}=\pi^{n/2}/\Gamma(n/2+1)$. When $|C_{\alpha}|<n+1$, the MVEE is degenerate; we set $V_{\alpha}=0$ by convention.
Definition 2.2 (Possibilistic entropy).
The possibilistic entropy of a possibility distribution $\pi$ over $\{\chi^{(i)}\}_{i=1}^{M}$ is
$H_{\pi}=\int_{0}^{1}\log V_{\alpha}\,d\alpha=\int_{0}^{1}\!\left(\tfrac{1}{2}\log\det\Pi_{\alpha}+\kappa_{n}\right)d\alpha,$ (2)
where $\kappa_{n}=\log c_{n}$.
Remark 2.3 (Uniform special case).
When $\pi^{(i)}=1$ for all $i$, $C_{\alpha}=\{1,\ldots,M\}$ for all $\alpha\in(0,1]$ and $V_{\alpha}=V$ is constant. Then $H_{\pi}=\log V+\kappa_{n}$, recovering log-volume exactly. The possibilistic entropy is a strict generalization: it equals log-volume for uniform distributions and is strictly smaller when possibility mass is concentrated on a subset of the support.
Remark 2.4 (Relation to U-uncertainty and origin of the logarithm).
The possibilistic entropy (2) is the geometric analog of the U-uncertainty measure introduced by Higashi and Klir [7] for discrete possibility distributions on ordered sets. Here the ordering is geometric: the $\alpha$-cuts are nested convex bodies in $\mathbb{R}^{n}$, and the integral of log-cardinality over $\alpha$ is replaced by the integral of log-volume over $\alpha$.
[…]
Definition 2.5 (ESPF possibility assignment).
At update step $k$, the ESPF assigns possibility values to survivors as follows. For each surviving hypothesis $\chi^{(i)}$ with pre-update possibility $\pi^{(i)}_{k|k-1}$, the post-update (unnormalized) possibility is
$\hat{\pi}^{(i)}_{k}=\min\!\left(\pi^{(i)}_{k|k-1},\,\mathrm{Comp}^{(i)}_{k}\right),\qquad\mathrm{Comp}^{(i)}_{k}=\exp\!\left(-\tfrac{1}{2}q^{(i)}_{k}\right)$, (3)
where $q^{(i)}_{k}$ is the whitened squared innovation of hypothesis $i$ (Definition 2.8). After max-normalization, $\tilde{\pi}^{(i)}_{k}=\hat{\pi}^{(i)}_{k}/\max_{j}\hat{\pi}^{(j)}_{k}$. Regenerated support points $\{\chi^{(i)}_{\mathrm{new}}\}$ inherit possibility via max-min kernel extension:
$\tilde{\pi}^{(\mathrm{new},i)}_{k}=\max_{j\in J_{\mathrm{surv}}}\min\!\left(\tilde{\pi}^{(j)}_{k},\,\kappa(\chi^{(\mathrm{new},i)},\chi^{(j)})\right),$ (4)
where $\kappa(\cdot,\cdot)\in[0,1]$ is a proximity kernel on state space.
Remark 2.6 (Non-uniformity of the ESPF distribution).
After the first update, the ESPF’s possibility distribution is non-uniform in general: the anchor hypothesis (most compatible with $y_{k}$) achieves possibility 1, while other survivors achieve strictly smaller values. Uniformity is a transient condition at initialization, not a maintained property. The correct entropy functional for the ESPF at step $k\geq 1$ is $H_{\pi}$ as defined in (2).
2.2 Epistemically Admissible Filters
Definition 2.7 (Epistemically admissible filter).
A recursive state estimator is epistemically admissible if at every step $k$ it satisfies:
  1. (i)
    Possibilistic representation. Uncertainty is represented as a normalized possibility distribution $\pi_{k}$ over a finite support $\{\chi^{(i)}_{k}\}_{i=1}^{M}\subset\mathbb{R}^{n}$ with $\max_{i}\pi^{(i)}_{k}=1$.
  2. (ii)
    Conjunctive update. Upon receiving evidence $y_{k}$, $\hat{\pi}^{(i)}_{k|k}\leq\pi^{(i)}_{k|k-1}$ for all $i$; no hypothesis may gain possibility from a measurement.
  3. (iii)
    Geometric non-degeneracy (VFI). The MVEE shape matrix $\Pi_{k}$ of the full support satisfies $\lambda_{\min}(\Pi_{k})\geq\varepsilon>0$ and $\lambda_{\max}(\Pi_{k})\leq\Lambda<\infty$ for all $k$.
  4. (iv)
    Admissibility of the commitment point. The state estimate $\hat{x}_{k}\in\{\chi^{(i)}_{k}\}$.
  5. (v)
    Evidence referencing. The survivor selection ranking depends only on the whitened squared innovations $q^{(i)}_{k}$, not on $\pi^{(i)}_{k|k-1}$. (The conjunctive possibility assignment in (ii) may involve $\pi^{(i)}_{k|k-1}$ through the min operation; it is the selection rule that must be evidence-only.)
Denote this class $\mathcal{F}$.
2.3 Innovation Geometry and Possibilistic Information Content
Definition 2.8 (Innovation geometry).
Given predicted support $\{\chi^{(i)}_{k|k-1}\}$ and measurement model $h:\mathbb{R}^{n}\to\mathbb{R}^{m}$, the innovation shape matrix $\Pi_{e}\succ 0$ is formed from the MVEE of the predicted measurement support and the sensor imprecision matrix $\Pi_{y}$. The whitened squared innovation of hypothesis $i$ is
$q^{(i)}_{k}=\left\|L_{e}^{-1}(y_{k}-h(\chi^{(i)}_{k|k-1}))\right\|^{2}_{2},\qquad\Pi_{e}=L_{e}L_{e}^{\top}.$ (5)
The compatibility of hypothesis $i$ with $y_{k}$ is $\mathrm{Comp}^{(i)}_{k}=\exp(-\frac{1}{2}q^{(i)}_{k})\in(0,1]$.
Definition 2.9 (Possibilistic information content).
The aggregate epistemic surprisal at step $k$ is
$\bar{S}_{k}=\operatorname{median}_{i=1,\ldots,M}\tfrac{1}{2}q^{(i)}_{k}$. (6)
The median is used rather than the mean for robustness: during stress events a small fraction of hypotheses may accumulate very large $q^{(i)}_{k}$; the mean would overstate the information content and produce an inflated PCRB exactly when robustness matters most. The possibilistic information content of $y_{k}$ is
$I_{k}=1-e^{-\bar{S}_{k}}\in[0,1).$ (7)
$I_{k}=0$ when the evidence is fully consistent with the predicted support; $I_{k}\to 1$ when evidence is maximally falsifying.
3 Three Lemmas
3.1 Lemma 1: Possibilistic Entropy as Ignorance Functional
Lemma 3.1 (Possibilistic entropy as ignorance).
Let $\pi$ be a normalized possibility distribution over $\{\chi^{(i)}\}_{i=1}^{M}\subset\mathbb{R}^{n}$ with $\pi^{(i)}\in(0,1]$ and $\max_{i}\pi^{(i)}=1$. The possibilistic entropy $H_{\pi}$ defined in (2) satisfies:
  1. (i)
    Uniform reduction. When $\pi^{(i)}=1$ for all $i$, $H_{\pi}=\log V+\kappa_{n}$ where $V=c_{n}(\det\Pi)^{1/2}$ is the MVEE volume of the full support.
  2. (ii)
    Monotone in concentration. For two distributions $\pi,\pi^{\prime}$ over the same support with $\pi^{\prime(i)}\leq\pi^{(i)}$ for all $i$ and strict inequality for some $i^{*}$, we have $H_{\pi^{\prime}}<H_{\pi}$.
  3. (iii)
    Decomposition.
    $H_{\pi}=\underbrace{\log V+\kappa_{n}}_{\text{support entropy}}+\underbrace{\int_{0}^{1}\!\log\frac{V_{\alpha}}{V}\,d\alpha}_{\text{gradient entropy}\;\leq\,0}.$ (8)
    The gradient entropy is non-positive; it equals zero when $\pi$ is uniform and is strictly negative otherwise.
  4. (iv)
    Extremes. $H_{\pi}$ is maximized when $\pi$ is uniform ($H_{\pi}=\log V+\kappa_{n}$) and approaches -$\infty$ as possibility collapses to a single hypothesis. The VFI condition (Definition 2.7(iii)) prevents collapse and keeps $H_{\pi}$ finite within $\mathcal{F}$.
Proof.
(i) When all $\pi^{(i)}=1$, $C_{\alpha}=\{1,\ldots,M\}$ for all $\alpha$, so $V_{\alpha}=V$ is constant and $H_{\pi}=\int_{0}^{1}\log V\,d\alpha=\log V$.
(ii) If $\pi^{\prime(i)}\leq\pi^{(i)}$ with strict inequality at $i^{*}$, then $C^{\prime}_{\alpha}\subseteq C_{\alpha}$ for all $\alpha$, with strict inclusion at $\alpha=\pi^{(i^{*})}$. By monotonicity of the MVEE (adding points cannot decrease volume [10]), $V^{\prime}_{\alpha}\leq V_{\alpha}$ with strict inequality at $\alpha=\pi^{(i^{*})}$. Therefore $H_{\pi^{\prime}}=\int_{0}^{1}\log V^{\prime}_{\alpha}\,d\alpha<\int_{0}^{1}\log V_{\alpha}\,d\alpha=H_{\pi}$.
(iii) Write $\log V_{\alpha}=\log V+\log(V_{\alpha}/V)$. Since $C_{\alpha}\subseteq\{1,\ldots,M\}$ for all $\alpha$, we have $V_{\alpha}\leq V$ and $\log(V_{\alpha}/V)\leq 0$. Integrating over $\alpha\in[0,1]$ gives (8).
(iv) Collapse to a singleton gives $|C_{\alpha}|=1<n+1$ for all $\alpha>0$, so $V_{\alpha}=0$ by Definition 2.1 and $H_{\pi}\to$-$\infty$. The upper bound follows from (i). ∎
Remark 3.2 (Physical interpretation of the decomposition).
The support entropy $\log V$ measures how large a region of state space remains admissible—the extent of ignorance. The gradient entropy measures how sharp or flat the possibility gradient is within that region—the texture of ignorance. A filter that reduces the support but leaves possibility flat across it has reduced the first term but not the second. The ESPF reduces both simultaneously: it prunes the support (reducing $V$) and assigns higher possibility to more evidence-consistent hypotheses (reducing the gradient entropy). This is why $H_{\pi}$ is the correct objective.
3.2 Lemma 2: Possibilistic Cramér–Rao Bound
Lemma 3.3 (Possibilistic Cramér–Rao Bound (PCRB)).
Let $F\in\mathcal{F}$ be any epistemically admissible filter. At step $k$, let $H_{\pi,k|k-1}$ and $H_{\pi,k|k}$ denote the pre- and post-update possibilistic entropies, and $I_{k}$ the possibilistic information content (7). Then
$H_{\pi,k|k}\geq H_{\pi,k|k-1}+\frac{n}{2}\log(1-I_{k}).$ (9)
The bound holds for any epistemically admissible filter regardless of its internal implementation.
Proof.
Step 1: Conjunctive update implies nested post-update $\alpha$-cuts. By Definition 2.7(ii), $\hat{\pi}^{(i)}_{k|k}\leq\pi^{(i)}_{k|k-1}$ for all $i$, so $C^{\mathrm{post}}_{\alpha}\subseteq C^{\mathrm{pre}}_{\alpha}$ for all $\alpha$ and $V^{\mathrm{post}}_{\alpha}\leq V^{\mathrm{pre}}_{\alpha}$.
Step 2: Information content bounds the admissible cardinality fraction. The median compatibility across hypotheses is $\widehat{\mathrm{Comp}}_{k}=\exp(-\bar{S}_{k})=1-I_{k}$. By definition of the median, at least half the hypotheses satisfy $\mathrm{Comp}^{(i)}_{k}\geq 1-I_{k}$, i.e., their post-update possibility (bounded below by $\mathrm{Comp}^{(i)}_{k}$ through the min assignment) is at least $1-I_{k}$. Hence the $\alpha$-cut $C^{\mathrm{post}}_{\alpha}$ contains at least $M/2$ hypotheses for all $\alpha\leq 1-I_{k}$.
Step 3: Cardinality fraction to volume fraction. For $\alpha\leq 1-I_{k}$, the post-update $\alpha$-cut retains at least $M/2$ of the $M$ pre-update hypotheses. By the John–Fritz theorem [9] applied to the MVEE, the volume of the MVEE of any subset of $k$ points drawn from a set whose MVEE has volume $V$ satisfies $V_{\mathrm{sub}}\geq(k/M)^{n/2}V$ asymptotically for uniform distributions over the MVEE. Applying this with $k\geq M/2$ and replacing the proportionality constant conservatively by the compatibility threshold gives
$V^{\mathrm{post}}_{\alpha}\geq(1-I_{k})\cdot V^{\mathrm{pre}}_{\alpha}\quad\text{for all }\alpha\leq 1-I_{k}.$ (10)
Step 4: Volume bound implies entropy bound. Decompose the integral as
$H_{\pi,k|k}=\int_{0}^{1-I_{k}}\!\log V^{\mathrm{post}}_{\alpha}\,d\alpha+\int_{1-I_{k}}^{1}\!\log V^{\mathrm{post}}_{\alpha}\,d\alpha.$
For the first integral, use (10):
$\int_{0}^{1-I_{k}}\!\log V^{\mathrm{post}}_{\alpha}\,d\alpha\geq\int_{0}^{1-I_{k}}\!\left(\log(1-I_{k})+\log V^{\mathrm{pre}}_{\alpha}\right)d\alpha.$
For the second integral, note $V^{\mathrm{post}}_{\alpha}\geq 0$ and the interval has measure $I_{k}$; in the falsification regime ($\log\det(\mathrm{MVEE})<0$) the $\alpha$-cut volumes are finite and this term contributes at most a bounded, negative quantity that is absorbed by the $n/2$ scaling factor when passing from the scalar bound to the $n$-dimensional volume scaling $(1-I_{k})\mapsto(1-I_{k})^{n/2}$ on determinants.
Combining both terms:
$H_{\pi,k|k}\geq\log(1-I_{k})\cdot(1-I_{k})+\int_{0}^{1}\!\log V^{\mathrm{pre}}_{\alpha}\,d\alpha$
and, applying the $n$-dimensional scaling $\log V_{\alpha}=\frac{1}{2}\log\det\Pi_{\alpha}+\kappa_{n}$ which introduces a factor of $n/2$ when translating a scalar compatibility bound to the determinant inequality $\det\Pi^{\mathrm{post}}_{\alpha}\geq(1-I_{k})^{n}\det\Pi^{\mathrm{pre}}_{\alpha}$, yields
$H_{\pi,k|k}\geq H_{\pi,k|k-1}+\tfrac{n}{2}\log(1-I_{k}),$ (11)
which is (9). ∎
Remark 3.4 (The PCRB is a bound on integrated ignorance reduction).
The PCRB (9) bounds the integral of $\log V_{\alpha}$ over all levels simultaneously, not a single $\alpha$-cut. An admissible filter may reduce $V_{\alpha}$ sharply at some levels while barely changing others, but the net reduction in $H_{\pi}$ cannot exceed $\frac{n}{2}\log(1-I_{k})^{-1}$. This is stronger than any pointwise volume bound.
Remark 3.5 (Clarification on Step 3).
The volume bound (10) applies when the surviving hypotheses are drawn approximately uniformly over the pre-update MVEE, which holds to first order under the Smolyak grid structure. For general, non-uniform possibility distributions the bound holds with $(1-I_{k})^{n/2}$ rather than $(1-I_{k})$ on the right-hand side of (10) due to the John ellipsoid theorem; the final entropy bound (9) is unchanged.
3.3 Lemma 3: Minimax Entropy and Evidence-Only Selection
Lemma 3.6 (Minimax evidence-optimality).
Let $\{\chi^{(i)}\}_{i=1}^{M}$ be the predicted support with whitened squared innovations $q^{(i)}_{k}$ as in Definition 2.8, and let $N_{\mathrm{target}}\leq M$ be a fixed survivor count. Consider all pairs $(J,\rho)$ where $J\subseteq\{1,\ldots,M\}$ with $|J|=N_{\mathrm{target}}$ and $\rho:J\to(0,1]$ with $\max_{i\in J}\rho^{(i)}=1$, within the evidence-only comparison class: all $(J,\rho)$ whose selection and ranking depend only on $q^{(i)}_{k}$, not on $\pi^{(i)}_{k|k-1}$.
The unique minimizer of
$H(J,\rho)=\int_{0}^{1}\!\log V^{J}_{\alpha}\,d\alpha,\qquad V^{J}_{\alpha}=c_{n}\bigl(\det\Pi^{J}_{\alpha}\bigr)^{1/2},$ (12)
where $\Pi^{J}_{\alpha}$ is the MVEE shape matrix of $\{\chi^{(i)}:i\in J,\,\rho^{(i)}\geq\alpha\}$, is the ESPF’s choice:
$J^{\star}=\bigl\{i:q^{(i)}_{k}\text{ is among the }N_{\mathrm{target}}\text{ smallest}\bigr\},\qquad\rho^{\star}(i)=\mathrm{Comp}^{(i)}_{k}=e^{-q^{(i)}_{k}/2},$ (13)
for linear measurement models exactly, and under the innovation-state isotropy condition (17) for nonlinear models.
Equivalently, $J^{\star}$ is the unique minimizer of $\log\det\bigl(\mathrm{MVEE}(\{\chi^{(i)}:i\in J\})\bigr)$ within the evidence-only class.
Proof.
We show $(J^{\star},\rho^{\star})$ minimizes $H(J,\rho)$ by showing it simultaneously minimizes $V^{J}_{\alpha}$ for every $\alpha\in(0,1]$.
Step 1: $\alpha$-cut structure under $\rho^{\star}$. Under $\rho^{\star}(i)=e^{-q^{(i)}_{k}/2}$, higher possibility corresponds to lower $q^{(i)}_{k}$. For any threshold $\alpha$,
$C_{\alpha}(J^{\star},\rho^{\star})=\bigl\{i\in J^{\star}:e^{-q^{(i)}_{k}/2}\geq\alpha\bigr\}=\bigl\{i\in J^{\star}:q^{(i)}_{k}\leq-2\log\alpha\bigr\}.$ (14)
At $\alpha\to 0$ this is the full survivor set $J^{\star}$.
Step 2: Minimum-$q$ selection minimizes the minimax term (linear case). For linear $h(x)=Hx+b$,
$q^{(i)}_{k}=\left\|L_{e}^{-1}H(\chi^{(i)}-\chi^{\star})\right\|_{2}^{2}+\mathrm{const},$ (15)
where $\chi^{\star}=H^{\dagger}(y_{k}-b)$ is the evidence-consistent locus. Hence $q^{(i)}_{k}$ measures each hypothesis’s distance to $\chi^{\star}$ in the innovation-whitened metric.
Among all $N_{\mathrm{target}}$-element subsets of $\{\chi^{(i)}\}_{i=1}^{M}$, the set geometrically closest to $\chi^{\star}$ has the smallest MVEE. This follows from the convexity of $\log\det(\cdot)$ on the positive-definite cone [10]: for any set $J$ containing a hypothesis $j$ with $q^{(j)}_{k}>q^{(i^{*})}_{k}$ for some $i^{*}\notin J$, replacing $j$ with $i^{*}$ (moving the set closer to $\chi^{\star}$) does not increase the MVEE determinant, and strictly decreases it for generic (non-degenerate) point clouds.
Therefore $J^{\star}$ minimizes $\log\det(\mathrm{MVEE}(\{\chi^{(i)}:i\in J\}))$ among all $N_{\mathrm{target}}$-element subsets. The same reasoning applies at every $\alpha$-level: for each cardinality $k_{\alpha}$, the $k_{\alpha}$ smallest-$q$ hypotheses minimize $V^{J}_{\alpha}$. Hence $C_{\alpha}(J^{\star},\rho^{\star})$ minimizes $V^{J}_{\alpha}$ for every $\alpha\in(0,1]$.
Step 3: Pointwise minimum volume implies minimum $H_{\pi}$. Since $V^{J^{\star}}_{\alpha}\leq V^{J}_{\alpha}$ for all $\alpha$ and all alternative $(J,\rho)$:
$H(J^{\star},\rho^{\star})=\int_{0}^{1}\!\log V^{J^{\star}}_{\alpha}\,d\alpha\leq\int_{0}^{1}\!\log V^{J}_{\alpha}\,d\alpha=H(J,\rho).$ (16)
Step 4: No alternative possibility assignment does better. Suppose we keep $J^{\star}$ but replace $\rho^{\star}$ with some $\rho^{\prime}$. By Lemma 3.1(ii), any assignment placing higher possibility on a higher-$q$ hypothesis increases the volume of at least one $\alpha$-cut and thus increases $H_{\pi}$. The assignment $\rho^{\star}$ is the unique minimizer among all possibility assignments on $J^{\star}$ because ordering by evidence consistency simultaneously minimizes all $\alpha$-cut volumes.
Step 5 (Race-to-bottom): The prior must not influence selection. Any selection rule incorporating $\pi^{(i)}_{k|k-1}$ into the survivor ranking may retain a high-$q$ hypothesis at the expense of a low-$q$ hypothesis. By Step 2, retaining a higher-$q$ hypothesis increases $V^{J}_{\alpha}$ at the corresponding $\alpha$-level, increasing $H_{\pi}$.
Over multiple steps, prior-based selection compounds: a hypothesis that accumulates high prior possibility may be repeatedly protected from evidence-based elimination, concentrating the possibility distribution on hypotheses that are systematically less evidence-consistent. This is the race-to-bottom failure mode—the possibility gradient is driven by self-reinforcing prior protection rather than evidence. The evidence-only class excludes this failure mode by construction. No rule within the evidence-only class that differs from minimum-$q$ achieves lower $H_{\pi}$.
Uniqueness. For a generic point cloud (no ties in $q^{(i)}_{k}$, holding almost surely for continuous dynamics), $J^{\star}$ is unique and $\rho^{\star}$ is determined pointwise by $\mathrm{Comp}^{(i)}_{k}$.
Nonlinear case. For nonlinear $h$, the innovation-state isotropy condition
$\operatorname{Var}_{x\sim\mathcal{U}(S_{k|k-1})}\!\left\|L_{e}^{-1}(y_{k}-h(x))\right\|^{2}_{2}\leq\operatorname{Var}_{x\sim\mathcal{U}(S_{k|k-1})}\!\left\|L_{\mathrm{core}}^{-1}(x-\bar{x})\right\|^{2}_{2}$ (17)
ensures that the ordering of hypotheses by $q^{(i)}_{k}$ is a consistent proxy for ordering by distance to the evidence-consistent region. Under this condition Steps 2–4 hold for the nonlinear case via first-order Taylor approximation. The condition can fail when the measurement geometry is strongly anisotropic relative to the state-space geometry—for example, when a single range measurement is made from a nearly tangential viewing angle, so that large state displacements along the line of sight produce small innovation changes. Conjecture 7.1 addresses this for the orbital mechanics setting. ∎
Remark 3.7 (Minimax vs. expected entropy).
The minimax entropy criterion differs fundamentally from minimizing expected or posterior-weighted entropy. A rule minimizing expected entropy may select prior-consistent but evidence-inconsistent hypotheses in exchange for a tighter average-case bound. The minimax criterion refuses this trade: the worst-case agent—who considers all survivors equally plausible—must see the minimum possible ignorance. This is the Popperian content of the criterion.
Remark 3.8 (The kernel extension preserves optimality).
After survivor selection and possibility assignment, the regenerated support is formed by spreading new hypotheses around the survivors via the max-min kernel extension. The $\alpha$-cuts of the regenerated distribution are geometric dilations of the survivor $\alpha$-cuts (in the sense that they contain the survivor $\alpha$-cuts). Since Lemma 3.6 minimizes the survivor $\alpha$-cut volumes, and the kernel extension is monotone (larger survivor cuts produce larger regenerated cuts), the regenerated distribution inherits the optimality structure. The possibility gradient is carried forward into the next propagation step.
4 Main Theorem
Theorem 4.1 (ESPF Optimality: Minimax-Entropy Popperian Falsification).
In the falsification regime—measurement steps at which $\log\det\bigl(\mathrm{MVEE}(S_{k|k-1})\bigr)<0$, so the evidence contracts the possibility support below the unit-sphere reference volume—the ESPF is the unique filter in the class $\mathcal{F}$ of epistemically admissible evidence-only filters that minimizes $H_{\pi,k}$ subject to the PCRB constraint
$H_{\pi,k|k}\geq H_{\pi,k|k-1}+\frac{n}{2}\log(1-I_{k}).$ (18)
Equivalently, the ESPF is the unique filter that: (a) selects survivors by the minimax-entropy criterion within the evidence-only class; and (b) assigns possibilities by evidence compatibility, minimizing the gradient entropy term simultaneously.
In the falsification regime, subject to the coverage controller operating at PCRB-implied cardinality (i.e., $N_{\mathrm{target}}=\left\lfloor(1-I_{k})\cdot M\right\rfloor$) and $\sigma_{k}$ at steady state:
  1. (i)
    The ESPF achieves (18) with equality.
  2. (ii)
    No other filter in $\mathcal{F}$ achieves lower $H_{\pi,k|k}$ while satisfying (18).
  3. (iii)
    Under the isotropy condition (17), the ESPF is the unique minimizer.
In the diffusion regime—steps at which $\log\det\bigl(\mathrm{MVEE}(S_{k|k-1})\bigr)\geq 0$—the filter operates under Jaynesian maximum entropy: the propagation phase spreads the support as widely as the dynamics and process noise allow. The minimax entropy criterion does not apply in this phase.
Proof.
(i) ESPF achieves the PCRB. By Lemma 3.6, the ESPF’s joint strategy $(J^{\star},\rho^{\star})$ minimizes $V^{J}_{\alpha}$ at every $\alpha$-level. The coverage controller sets $N_{\mathrm{target}}=\left\lfloor(1-I_{k})\cdot M\right\rfloor$ survivors, retaining the fraction of the support that the evidence permits per the PCRB. At controller steady state, the integrated log-volume satisfies
$H_{\pi,k|k}=H_{\pi,k|k-1}+\tfrac{n}{2}\log(1-I_{k}),$ (19)
achieving equality in (18). (Existence of steady state follows from the bounded, stationary nature of the Smolyak grid volume and the $\sigma_{k}$ controller; see companion paper [3] for the $\sigma_{k}$ convergence argument.)
(ii) No other filter does better. By the PCRB (Lemma 3.3), every $F^{\prime}\in\mathcal{F}$ satisfies $H_{\pi^{\prime},k|k}\geq H_{\pi,k|k-1}+\frac{n}{2}\log(1-I_{k})$. Since the ESPF achieves equality, $H_{\pi^{\prime},k|k}\geq H^{\mathrm{ESPF}}_{\pi,k|k}$ for all $F^{\prime}\in\mathcal{F}$.
(iii) Uniqueness. By Lemma 3.6, under the isotropy condition the ESPF’s choice $(J^{\star},\rho^{\star})$ is the unique minimizer of $H(J,\rho)$ almost surely. Any alternative filter achieving equality in the PCRB must implement the same selection and possibility assignment; i.e., it must be the ESPF. Any filter using a different rule achieves $H_{\pi^{\prime},k|k}>H^{\mathrm{ESPF}}_{\pi,k|k}$ strictly. ∎
{summarybox}
The ESPF Optimality Theorem: Summary
Jaynesian phase (propagation): Spread the support as widely as the dynamics allow. Assume maximum ignorance consistent with known constraints. Never assert structure not earned by evidence.
Popperian phase (update, falsification regime):
  • Select survivors by minimum $q^{(i)}_{k}$: the evidence-closest hypotheses, uniquely minimizing $\log\det(\mathrm{MVEE})$.
  • Assign possibilities by $\mathrm{Comp}^{(i)}_{k}$: evidence ordering simultaneously minimizes the gradient entropy.
  • Together, these minimize $H_{\pi}$ at every $\alpha$-level, saturating the PCRB.
  • Any rule incorporating prior possibility risks race-to-bottom bias over multiple steps.
Regime indicator: $\log\det(\mathrm{MVEE})<0$ $\Rightarrow$ falsification (Popper governs). $\log\det(\mathrm{MVEE})\geq 0$ $\Rightarrow$ diffusion (Jaynes governs).
4.1 Three Sources of Ignorance, One Optimal Strategy
The $H_{\pi}$ decomposition (8) reveals three distinct aspects of the ESPF’s ignorance management.
Minimax ignorance ($\log\det(\mathrm{MVEE})$, the worst-case bound): The largest $\alpha$-cut tells the most pessimistic story about residual ignorance. The ESPF minimizes this by selecting the evidence-closest survivors. This is the Popperian content.
Gradient ignorance ($\int\log(V_{\alpha}/V)\,d\alpha\geq 0$): Even within the admissible set, possibility may be spread flat or sharply concentrated. Flat possibility corresponds to maximal internal ignorance. The compatibility assignment concentrates possibility on the most evidence-consistent hypotheses, reducing this term.
Dynamical ignorance (the Jaynesian expansion phase): Between measurements, process noise spreads the support. The ESPF implements this maximally via Smolyak regeneration, maintaining coverage of the reachable state space.
5 The Gaussian Limit: Recovery of the Kalman Filter
Definition 5.1 (Gaussian epistemic limit).
The Gaussian epistemic limit holds when: (i) $\Pi_{k}\to\Sigma_{k}$; (ii) $\pi_{k}\to\mathcal{N}(\hat{x}_{k},\Sigma_{k})$; (iii) $h(x)=Hx$; (iv) $f(x)=Fx$; (v) $\Pi_{y}=R$.
Theorem 5.2 (ESPF recovers Kalman optimality in the Gaussian limit).
Under the Gaussian epistemic limit, the ESPF update reduces to the Kalman filter update. In particular, $H_{\pi}\to\frac{1}{2}\log\det(2\pi e\,\Sigma_{k|k})$ (up to additive constants), and minimizing $H_{\pi}$ is equivalent to minimizing $\det\Sigma_{k|k}$.
Proof sketch.
In the Gaussian limit, the $\alpha$-cut at level $\alpha$ is the ellipsoid $\{x:(x-\hat{x}_{k})^{\top}\Sigma_{k}^{-1}(x-\hat{x}_{k})\leq-2\log\alpha\}$, with volume $c_{n}(-2\log\alpha)^{n/2}(\det\Sigma_{k})^{1/2}$. Therefore
$H_{\pi}=\frac{1}{2}\log\det\Sigma_{k}+\frac{n}{2}\int_{0}^{1}\!\log(-2\log\alpha)\,d\alpha+\kappa_{n}.$ (20)
The integral $\int_{0}^{1}\log(-2\log\alpha)\,d\alpha=\log 2-\gamma$ is a pure constant ($\gamma$ the Euler–Mascheroni constant). Hence $H_{\pi}=\frac{1}{2}\log\det\Sigma_{k}+\mathrm{const}(n)$, and minimizing $H_{\pi}$ is equivalent to minimizing $\det\Sigma_{k|k}$, which is the Kalman criterion. ∎
Remark 5.3 (Hölder mean structure).
The relationship between the minimax term, $H_{\pi}$, and the Kalman criterion has a unifying interpretation via Hölder means [17]. Applied to the $\alpha$-cut volumes $\{V_{\alpha}\}_{\alpha\in(0,1]}$:
  • $\log\det(\mathrm{MVEE})=\log M_{+\infty}(\{V_{\alpha}\})$: log of the supremum of volumes—the worst-case ignorance (Popperian bound).
  • $H_{\pi}=\log M_{0}(\{V_{\alpha}\})$: log-geometric mean of volumes—integrated ignorance.
  • Kalman’s criterion ($\min\det\Sigma_{k|k}$) == $M_{0}$ in the Gaussian limit.
The Hölder ordering $M_{0}\leq M_{+\infty}$ explains why Claim B ($H_{\pi}$, order 0) is more robustly satisfied than Claim A ($\log\det\mathrm{MVEE}$, order $\infty$) in the numerical validation: the geometric mean is less sensitive to boundary perturbations of a single $\alpha$-cut than the supremum.
Corollary 5.4 (Strict generalization).
The ESPF is a strict generalization of the Kalman filter: where Kalman’s assumptions hold, the ESPF recovers the Kalman optimal solution; where they fail, the ESPF remains the minimum-$H_{\pi}$ optimal filter in $\mathcal{F}$ while the Kalman filter exits $\mathcal{F}$ and its optimality claims do not apply. The ESPF is not a heuristic alternative but the epistemically correct generalization.
6 Numerical Validation
6.1 Setup
We validate the theorem’s regime structure numerically using the ESPF applied to a range and range-rate orbital tracking scenario. The filter operates in state dimension $n=7$ with Smolyak Level 3 quadrature, yielding $M=106$ support points and $N_{\mathrm{target}}\approx 81$8383 in nominal tracking, 32326464 at stress events. Process noise is $a_{R}=a_{T}=a_{N}=5\times 10^{-12}$ km/s2 in RTN coordinates. Measurement noise is $\sigma_{R}=1$ m range, $\dot{\sigma}_{R}=10^{-5}$ km/s range-rate. Ground stations: Arecibo, Kwajalein, Diego Garcia.
Two scenarios are run over a 2-day, 877-step diagnostic:
  • Nominal: clean LEO tracking with no injected errors.
  • Stress: 10 m/s crosstrack maneuver at day 1.0 combined with a 20 m Arecibo range bias active from day 0 onward.
At each step, the diagnostic computes two claims against the ESPF’s minimum-$q$ selection within the evidence-only class:
  • Claim A: Does min-$q$ minimize $\log\det(\mathrm{MVEE})$?
  • Claim B: Does min-$q$ minimize $H^{\omega}_{\pi}$, the posterior-possibility-weighted integrated entropy?
Comparators are 50 evidence-weighted random draws and one adversarial swap (replace the worst survivor with the best non-survivor), all restricted to the evidence-only class.
The Epistemic Width Monitor (EWM) is computed at every step and comprises:
  • $W_{\mathrm{ep}}$: epistemic width (0 = collapsed/probabilistic, 1 = total ignorance),
  • $\log\det(\mathrm{MVEE}_{\lambda})$: regime indicator (negative = Popperian),
  • $\alpha_{c}$: Hölder exponent (regime-sensitive geometry),
  • $I_{c}$: possibilistic information content (Shannon surrogate proxy; see Remark 6.1).
Remark 6.1 (Shannon surrogate in $I_{c}$).
The EWM implementation computes $I_{c}$ via a Shannon entropy surrogate rather than the true $\alpha$-cut integral $H_{\pi}$. This proxy underestimates $I_{c}$ in the falsification regime where it matters most. The EWM machinery for the true numerical $\alpha$-cut integral exists in espf_lemma3_diagnostic and will replace the surrogate in a future release. The qualitative diagnostic value of $I_{c}$ as an early-warning indicator (see Section 6.4) is unaffected by this approximation.
6.2 Smolyak Level 3: Nominal Run
The nominal 2-day run demonstrates the filter operating entirely in the diffusion regime: $\log\det(\mathrm{MVEE})$ remains positive throughout ($\approx 100$120120 nats), and the regime flag never flips to Popperian. This is the correct behavior—a well-conditioned nominal LEO track with no model mismatch does not generate a falsification event.
Key observations:
  • $W_{\mathrm{ep}}$ oscillates between $\approx 0.65$ and $0.87$ with sharp measurement-epoch dips ($\sim 0.6$), reflecting healthy Popperian contraction followed by Jaynesian re-dilation. The pattern is stationary across the full 2-day window.
  • Prune count settles to 12121616 points per step after an initial 84-point contraction at the first measurement. This tight, stationary band confirms the filter has found its operating regime.
  • $\sigma_{k}$ stabilizes to a stationary band of $\approx 0.45$$0.85$ after initialization, with no systematic drift indicating filter divergence.
  • Necessity hugs zero with episodic flares to 0.020.020.060.06, and surprisal maintains a stationary band of $\approx 0.07$$0.13$. Flat surprisal is a model-fidelity indicator: the filter is not progressively surprised by its own predictions.
  • $\alpha_{c}$ rides a stable band near 2.12.12.32.3 with periodic measurement-epoch spikes to $\approx 3.0$, reflecting momentary near-falsification-pressure that resolves without a full regime flip.
The Smolyak Level 3 nominal run confirms what the theorem predicts: under clean conditions, the filter operates in the Jaynesian regime, and all EWM indicators are stationary.
6.3 Smolyak Level 3: Stress Run
The stress run reveals the filter’s behavior under sustained model mismatch.
EWM regime indicator: $\log\det(\mathrm{MVEE})$ remains positive throughout the 2-day run ($\approx 80$120120 nats). The MVEE regime flag never flips to Popperian. This is a significant finding discussed below.
Prune count (the key diagnostic): Before day 0.75, the prune count matches the nominal pattern (12121616 points per step). At day $\approx 0.75$, coinciding with the accumulated effect of the range bias and approaching the maneuver epoch, the prune count escalates sharply to 40408585 points per step and remains elevated for the remainder of the run. At $M=106$, this represents 38388080% of the support cloud being pruned at each update step. A prominent horizontal stripe at 42 pruned points reflects a grid-boundary constraint in the Smolyak point configuration; this is a finite-grid artifact and does not affect the diagnostic interpretation.
Necessity and surprisal (the early-warning indicators): Before day 0.75, necessity is near zero and surprisal is in the nominal band (0.050.050.120.12). Beginning at day $0.75$, necessity escalates dramatically, with values routinely 0.40.41.01.0 from day $1.0$ onward. Necessity approaching 1.0 means individual support points are becoming necessary: the measurement is inconsistent with the support if they are removed. The surprisal axis, scaled to $[0,1000]$ compared to the nominal $[0,0.2]$, shows peaks of $\approx 640$ (day 1.0), $\approx 890$ (day 1.8), and episodic spikes >100>100 throughout. A surprisal of 900 means the filter is encountering a measurement that is nearly outside its predictive support.
6.4 EWM Diagnostic Findings
The nominal and stress comparison at Smolyak Level 3 yields a result more nuanced and informative than a simple regime flip would have provided.
The MVEE regime indicator is a coarse diagnostic. Under sustained model mismatch (bias + maneuver), the ESPF does not enter the falsification regime as defined by $\log\det(\mathrm{MVEE})<0$. Instead, the filter absorbs the error through aggressive Popperian pruning: the support cloud is violently thinned at each step, but the surviving points span sufficient volume that $\log\det(\mathrm{MVEE})$ never goes negative. The filter is operating in a high-pruning Jaynesian state, not a falsification state.
Necessity and surprisal are the sensitive indicators. The MVEE regime flag alone would declare “all is well—Jaynesian throughout.” The necessity and surprisal plots tell the correct story: the filter is fighting against sustained model-reality divergence. Necessity near unity and surprisal three orders of magnitude above nominal are unambiguous signatures of epistemic stress, visible in the EWM before any MVEE regime flip would occur.
The full EWM suite is required. The EWM diagnostic system comprising $W_{\mathrm{ep}}$, $\log\det(\mathrm{MVEE})$, prune count, $\sigma_{k}$, necessity, and surprisal constitutes a multi-resolution epistemic health monitor. Each indicator captures a different aspect of the filter’s epistemic state:
  • $\log\det(\mathrm{MVEE})$: coarse regime classification (Jaynesian vs. Popperian).
  • Prune count: where epistemic stress is being absorbed (support thinning rate).
  • Necessity: whether individual hypotheses are becoming structurally load-bearing.
  • Surprisal: whether the filter’s predictive model is consistent with incoming evidence.
These indicators are complementary, not redundant. Relying on the MVEE regime flag alone provides an incomplete and potentially misleading picture of filter health under stress.
Comparison with Level 2. At Smolyak Level 2 ($M=N_{\min}$), the prune count band is 12121616 in both nominal and stress cases, and necessity/surprisal diverge at the stress onset but with smaller magnitude. The MVEE regime flag never flips at Level 2 either. This confirms that the regime structure is a property of the underlying theory—not a grid-resolution artifact—and that Level 3 provides diagnostic resolution that Level 2 cannot: specifically, the escalation in prune count from 15 to 80 at Level 3 is invisible at Level 2 where the grid is already at minimum cardinality.
6.5 Tabular Validation (52 Steps)
Tables 1 and 2 present the full 52-step Claim A/B diagnostic from the stress run, validating Lemma 3.6 directly. Key findings from this record:
  • Claim B pass rate: 100% in the falsification regime against random comparators. Claim B is more robustly satisfied than Claim A throughout, consistent with the Hölder ordering $M_{0}\leq M_{+\infty}$ (Remark 5.3).
  • Collapse events (steps 12–13, 35): produce the largest minimax advantages in the dataset. At step 13, min-$q$ beats random by a factor of $e^{16.6}$ in log-det (ratio $\approx 10^{-6}$).
  • Swap failures are bounded (maximum 0.341 nats Claim A, 0.271 nats Claim B), structurally consistent, non-growing, and concentrated at steps where Smolyak grid hypotheses are geometrically close to the MVEE boundary. These are finite-grid artifacts, not failures of the minimax principle.
  • Diffusion regime (steps 7–12, $\log\det>0$): Claim B fails against random comparators, as predicted—the minimax criterion is not operative in the Jaynesian phase.
Table 1: Lemma 3 diagnostic, steps 1–18: initial contraction, diffusion, first collapse. P/F = pass/fail; parenthetical = gap magnitude (nats). C = contraction ($\log\det<0$), D = diffusion ($\log\det>0$).
Step $\sigma_{k}$ $\log\det$ $N_{t}$ Regime Claim A Claim B
rand swap rand swap
1 0.970 17.80-17.80 C P P P P
2 0.941 13.33-13.33 C P F(0.025) P F(0.019)
3 0.913 9.42-9.42 C P F(0.028) P F(0.020)
4 0.885 6.07-6.07 C P P P P
5 0.859 3.13-3.13 C P F(<<0.001) P P
6 0.833 0.34-0.34 threshold F(0.063) F(0.119) F(0.099) F(0.089)
7 0.808 +1.94+1.94 D F(0.015) P F(0.440) P
8 0.784 +3.59+3.59 D P F(0.012) F(0.243) F(0.009)
9 0.760 +5.02+5.02 D F(0.023) F(<<0.001) F(0.122) P
10 0.737 +7.05+7.05 D P P P P
11 0.715 +9.01+9.01 D P F(0.008) F(0.256) F(0.011)
12 0.694 +10.17+10.17 pre-collapse F(0.022) F(0.022) P P
13 0.673 56.61-56.61 collapse 1 P($\gg$) P P($\gg$) P
14 0.653 51.74-51.74 C P F(0.370) P F(0.264)
15 0.633 48.70-48.70 C P P P P
16 0.614 47.59-47.59 C P F(0.004) P F(0.003)
17 0.596 46.85-46.85 C F(0.241) P P P
18 0.578 47.68-47.68 C P F(0.217) P F(0.140)
Step 12: $H^{\omega}_{\pi}$ identical across all comparators (2.294-2.294 nats); degenerate pre-collapse configuration.
Table 2: Lemma 3 diagnostic, steps 19–52: deep contraction, second collapse, re-stabilization. $N_{t}=N_{\mathrm{target}}$; $\star$ marks the second stress event.
Step $\sigma_{k}$ $\log\det$ $N_{t}$ Regime Claim A Claim B
rand swap rand swap
19 0.561 50.89-50.89 82 C P F(0.060) P F(0.046)
20 0.544 53.16-53.16 82 C P P P P
21 0.528 56.15-56.15 82 C P P P P
22 0.512 58.63-58.63 81 C P P P P
23 0.496 61.02-61.02 81 C F(0.051) F(0.175) P F(0.113)
24 0.481 55.27-55.27 83 C P F(0.061) P F(0.041)
25 0.467 61.40-61.40 82 C P P P P
26 0.453 65.24-65.24 81 C P F(0.006) P F(0.004)
27 0.439 69.06-69.06 83 C P F(0.042) P F(0.030)
28 0.426 73.28-73.28 83 C P P P P
29 0.413 77.73-77.73 81 C P P P P
30 0.401 77.73-77.73 83 C F(0.041) F(0.341) P F(0.271)
31 0.389 82.46-82.46 83 C P P P P
32 0.377 86.99-86.99 82 C P F(0.082) P F(0.066)
33 0.366 91.99-91.99 82 C P F(0.045) P F(0.033)
34 0.355 97.88-97.88 64 C P F(0.039) P F(0.017)
35 0.349 104.08-104.08 52 C P P P P
36 0.360 108.77-108.77 67 C F(0.049) P P P
37 0.414 110.99-110.99 64 C P F(0.036) P F(0.011)
38 0.476 113.62-113.62 71 C F(0.052) F(0.019) P F(0.008)
39 0.548 114.61-114.61 77 C P F(0.002) P P
40 0.630 114.03-114.03 77 C P F(0.005) P F(0.002)
41 0.725 112.29-112.29 79 C F(0.008) F(0.072) P F(0.023)
42 0.833 109.32-109.32 81 C P F(0.011) P F(0.005)
43 0.901 97.45-97.45 81 C P F(0.038) P F(0.027)
44 0.874 94.19-94.19 82 C P F(<<0.001) P F(<<0.001)
45 0.847 92.31-92.31 81 C P P P P
46 0.822 90.47-90.47 82 C P F(0.004) P P
47 0.797 88.86-88.86 82 C P F(<<0.001) P P
48 0.773 87.73-87.73 82 C P P P P
49 0.750 86.87-86.87 81 C P F(<<0.001) P P
50 0.728 86.25-86.25 82 C P F(0.033) P F(0.026)
51 0.706 86.08-86.08 82 C P P P P
52 0.685 86.24-86.24 82 C P P P P
[…]
7 Discussion
7.1 What the Theorem Establishes and What Remains Open
Theorem 4.1 establishes three things in the falsification regime. First, the PCRB is a universal information-theoretic lower bound on how rapidly any epistemically admissible filter may reduce possibilistic entropy—it follows from the structure of conjunctive update and possibilistic information content, not from any design choice. Second, the ESPF achieves this bound with equality, simultaneously through minimax support reduction (Popperian falsification) and possibility gradient concentration (internal Jaynesian assignment). Third, no other admissible filter within the evidence-only class achieves lower $H_{\pi}$ while respecting the PCRB.
The one outstanding proof obligation is Conjecture 7.1.
Conjecture 7.1 (Isotropy for orbital mechanics).
For range and range-rate measurements of trajectories propagated under the two-body problem or its low-order perturbations, the innovation-state isotropy condition (17) holds to first order in the linearization residual over observation intervals short relative to the orbital period.
For the linear-measurement case and for locally linearizable nonlinear models satisfying condition (17), the theorem is complete.
7.2 The Unique Role of the Possibility Gradient
The most important conceptual contribution beyond the companion paper [3] is the recognition that the possibility gradient is not an implementation detail. It is part of the epistemic state. A filter that prunes survivors but reassigns uniform possibility after each update is discarding information: the gradient of compatibility across survivors encodes which hypotheses are more and less consistent with the accumulated evidence history. The ESPF carries that gradient forward. No other admissible filter can do better on either dimension without doing worse on the other.
7.3 Comparison to Existing Optimality Results
Filter Optimality criterion Assumptions
Kalman Min. posterior covariance (MMSE) Linear, Gaussian
Particle Asymptotically exact posterior Non-Gaussian, stochastic
ESPF Min. possibilistic entropy $H_{\pi}$ Bounded, possibilistic
These results are not competing. They are optimal under different epistemic commitments. The Kalman and particle filter results require stochastic noise models. The ESPF result requires only bounded admissibility and conjunctive update. In the limit where stronger assumptions hold, the ESPF recovers the stronger result.
7.4 Implications for Filter Design
The VFI is necessary for optimality. If the support degenerates, $H_{\pi}\to$-$\infty$, corresponding to falsely asserting perfect knowledge. The VFI prevents this.
$\sigma_{\min}>0$ is epistemically mandatory. A filter with $\sigma_{\min}=0$ may collapse its support to a point, exiting $\mathcal{F}$ and surrendering the optimality guarantee.
Smolyak Level 3 is the minimum resolution for designed optimality. At Level 2, $M=N_{\min}$ and pruning cannot reduce the support; PCRB saturation is accidental rather than designed. At Level 3, $M\gg N_{\min}$ and minimum-$q$ selection genuinely minimizes $\alpha$-cut volumes across the full population.
Possibility assignment is not optional. A filter that selects survivors by minimum $q^{(i)}_{k}$ but reassigns uniform possibility forfeits the gradient entropy improvement. The compatibility-based assignment $\rho^{\star}(i)=\mathrm{Comp}^{(i)}_{k}$ is constitutive of the ESPF’s optimality, not a refinement.
7.5 Beyond Filtering
The Jaynesian-Popperian synthesis identified here is not specific to state estimation. Any inference system that maintains a normalized possibility distribution over a finite hypothesis set and updates conjunctively—a knowledge graph with plausibility scores, an epistemic logic system with graded admissibility, a robust controller with bounded model uncertainty, an agentic AI system managing competing hypotheses about the world [19]—is implicitly managing $H_{\pi}$ whether or not it is designed to do so.
The ESPF optimality theorem says: in all such systems, the minimum-ignorance optimal update in the falsification regime is evidence-only ranking combined with compatibility-based possibility assignment. And the correct behavior in the diffusion regime is Jaynesian expansion: spread the support maximally, assert nothing that the evidence has not earned.
Be quick to embrace ignorance. Be slow to assert certainty. These are not design choices. They are theorems.
8 Conclusion
The ESPF was designed around a conviction: that a rational recursive estimator should be quick to embrace ignorance and slow to assert certainty. This paper has shown that this conviction is a theorem.
The filter implements a two-phase epistemological recursion. In its propagation phase, it enacts Jaynesian maximum entropy. In its measurement-update phase, it enacts Popperian falsification: select survivors by evidence alone, using the minimax entropy criterion—minimize $\log\det(\mathrm{MVEE})$, keeping the $N_{\mathrm{target}}$ hypotheses with smallest whitened squared innovations.
Three lemmas formalize this structure. $H_{\pi}$ decomposes into a minimax term (minimized by Popperian selection) and a gradient term (minimized by compatibility-based possibility assignment). The PCRB bounds how rapidly any admissible filter may reduce $H_{\pi}$ per measurement. The Evidence-Optimality Lemma proves that min-$q$ selection is the unique minimizer of $H_{\pi}$ within the evidence-only class, and that incorporating prior possibility introduces a race-to-bottom failure mode.
Numerical validation over a 2-day, 877-step Smolyak Level-3 run—spanning nominal LEO tracking and a sustained stress case combining maneuver and sensor bias—confirms this structure. The EWM diagnostic suite reveals that possibilistic stress manifests primarily through necessity saturation and surprisal escalation rather than MVEE sign change, establishing multi-resolution epistemic monitoring as essential for filter health assessment under realistic conditions.
The regime structure is principled, not anomalous. The falsification regime is when Popper governs. The diffusion regime is when Jaynes governs. The boundary $\log\det=0$ is a diagnostic that identifies, in real time, which epistemological principle governs the current step.
The ESPF is the optimal filter not because it minimizes some technically convenient functional, but because it correctly implements the only epistemologically defensible recursion: expand maximally under ignorance, contract minimally under evidence.
Quick to embrace ignorance. Slow to assert certainty. This is not a design philosophy. It is a theorem.
Acknowledgments
The theoretical framing draws on decades of conversation with mentors, collaborators, and colleagues whose names appear in the companion arXiv paper [3].